What is the absorbed dose when 3 mJ of energy is absorbed by a mass of 10 kg?


1
Expert's answer
2021-02-25T11:28:14-0500

By definition, the absorbed dose is measure of the energy deposited in matter by ionizing radiation per unit mass:


"D = \dfrac{E}{m}"

where "E = 3mJ = 3\times 10^{-3}J" is the total absorbed energy, and "m = 10kg" is the mass. Thus, obtain:


"D = \dfrac{3\times 10^{-3}J}{10kg} = 3\times 10^{-4}Gy"

Answer. "3\times 10^{-4}Gy".